OpenStudy (anonymous):

Compare and Contrast: Below are two expressions. Simplify each and then choose the statement that is true. Expression #1 Expression #2 (3x^2)3x^2 (3x^3)^2(x^2) A. The exponents in Expression #1 are greater than the exponents of Expression #2. B. The exponents on Expression #2 are greater than the exponents of Expression #1. C. The exponents of Expression #1 are the same as the exponents of Expression #2. D. The relationship cannot be determined with the given information.

OpenStudy (anonymous):

\[3x^2\times 3x^2=9x^4\] wheras \[(3^2)^2\times x^2=9x^4\times x^2=9x^6\]

OpenStudy (anonymous):

So its B im sorry i second guess myself

OpenStudy (anonymous):

i would go with B. The exponents on Expression #2 are greater than the exponents of Expression #1. yes

OpenStudy (anonymous):

can i ask u a question i got the answer too but im not to sure

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

Choose the correct simplification of the expression (3x)4

OpenStudy (anonymous):

i got 12x^4

OpenStudy (anonymous):

is it \((3x)^4\) ?

OpenStudy (anonymous):

Yes sorry bout that

OpenStudy (anonymous):

no problem, just making sure but your answer is not quite right

OpenStudy (anonymous):

\[(3x)^4=3^4x^4=81x^4\]

OpenStudy (anonymous):

you do not multiply 3 times 4 raise 3 to the fourth power

OpenStudy (anonymous):

ohh ok i see what i did wrong thank you for the help :)

OpenStudy (anonymous):

yw